# Edit As You Wish: Video Caption Editing with Multi-grained User Control

### v2

@@ Abstract.
Automatically narrating videos in natural language complying with user requests, i.e. Controllable Video Captioning task, can help people manage massive videos with desired intentions. However, existing works suffer from two shortcomings: 1) the control signal is single-grained which can not satisfy diverse user intentions; 2) the video description is generated in a single round which can not be further edited to meet dynamic needs. In this paper, we propose a novel Video Caption Editing (VCE) task to automatically revise an existing video description guided by multi-grained user requests. Inspired by human writing-revision habits, we design the user command as a pivotal triplet {operation, position, attribute} to cover diverse user needs from coarse-grained to fine-grained. To facilitate the VCE task, we automatically construct an open-domain benchmark dataset named VATEX-EDIT and manually collect an e-commerce dataset called EMMAD-EDIT. We further propose a specialized small-scale model (i.e., OPA) compared with two generalist Large Multi-modal Models to perform an exhaustive analysis of the novel task. For evaluation, we adopt comprehensive metrics considering caption fluency, command-caption consistency, and video-caption alignment. Experiments reveal the task challenges of fine-grained multi-modal semantics understanding and processing. Our datasets, codes, and evaluation tools are ready to be open-sourced.
@@ 2. Related Work
Image Caption Editing. Conventional image captioning (Vinyals et al., 2015; Xu et al., 2015; Liu et al., 2016; Stefanini et al., 2022; Li et al., 2016; Liu et al., 2018; Yao et al., 2023; Chen et al., 2017; Yang et al., 2023a) generates the description for images from scratch which may lead to factual mistakes. Sammani and Elsayed (2019) firstly define the image caption editing task that modifies an existing caption (a.k.a. reference caption) conditioned on the image content to obtain more accurate descriptions. Sammani and Melas-Kyriazi (2020) further propose a novel EditNet framework to achieve interactive and adaptive edits. Yuan et al. (2023) design an adaptive text-denoising network to alleviate the semantic gap between input images and reference sentences. The above works all edit image captions implicitly. Wang et al. (2022) propose the explicit image caption editing task to make the modification process more explainable and efficient. In summary, these caption editing works can only correct the wrong content in the reference captions and ignore specific edit intentions of different users. In this paper, we integrate multi-grained user commands into the video description editing process.
Large Multimodal Models. Recent months have witnessed the tremendous success of Large Language Models (LLMs) (Chowdhery et al., 2022; Ouyang et al., 2022; Zeng et al., 2022; Touvron et al., 2023b, a; OpenAI, 2022, 2023a) towards artificial general intelligence. Further, Large Multimodal Models (LMMs) (Dai et al., 2023; Liu et al., 2023; Bai et al., 2023; Ye et al., 2023a; Cha et al., 2023; Li et al., 2023a; Zhang et al., 2023; Luo et al., 2023; Lin et al., 2023; Ren et al., 2023) endow LLMs with the visual understanding ability by incorporating vision backbones (Sun et al., 2023; Radford et al., 2021; Li et al., 2022b). Existing LMMs primarily bifurcate into two categories: Image Large Language Models (ImgLLMs) (Dai et al., 2023; Liu et al., 2023; Bai et al., 2023; Ye et al., 2023a; Cha et al., 2023) and Video Large Language Models (VidLLMs) (Li et al., 2023a; Zhang et al., 2023; Lin et al., 2023; Luo et al., 2023; Ren et al., 2023). The standard architecture of an LMM comprises a vision backbone for encoding images or videos, a projector (Li et al., 2023b; Cha et al., 2023) to condense and translate visual embeddings into textual semantic space, and an LLM to process all multimodal contexts. As the VCE task involves capturing video semantics, understanding textual user controls, and enabling multi-grained text editing, it can serve as a new touchstone for LMMs.
@@ 3.2. User Edit Command
It is not trivial to define flexible edit commands in the VCE task to meet various realistic user needs. We observe that natural language and writing-revision traces are two natural interactive modes. The former can be received from keyboards or speech converters, while the latter conveniently expresses user intentions with the prevalence of tablets and wireless stylus pens111https://www.apple.com/apple-pencil/. A command representation compatible with the above two signals is important and meaningful. In this paper, we propose a novel command representation in a triplet format {operation, position, attribute}, where operations control the overall description editing, positions specify the editing locations, which could affect the syntax of sentences, and attributes guide the editing operation to control the semantic contents of descriptions.
We define the atomic edit operations as add and delete, considering that the replace editing can be decomposed into the two atomic operations (i.e. first delete then add). Meanwhile, position and attribute in the triplet are optional, therefore, as shown in Table 1, seven specific commands222Note that we omit the command “del, pos, attr, delete attributes at specified positions”, as it can be covered by “del, pos, -, delete description at specified positions”. via different combinations of operation, position, and attribute elements in the triplet can cover multi-grained realistic demands from coarse-grained (global) controls to fine-grained (local) controls. The designed triplet command can be flexibly obtained by processing the inputs from front-end interfaces including natural language and writing-revision traces (details in Appendix E). In the following method and experiments sections, we perform video description editing directly based on the triplet command.
@@ 4.1. VATEX-EDIT Construction
We sample an annotated caption of a video as the reference caption, and the next goal is to construct the command and the related edited caption. In general, we aim to construct related (command, edited caption) samples including: 1) coarse-grained length-control commands referring to the global add or delete edits that result in length changes, and 2) finer-grained attribute-related commands to achieve add or delete attributes on the reference caption.
Coarse-grained length-control commands. For add operation, we directly select a longer caption with significant length differences from the rest annotations as the edited caption. The delete operation does the exact opposite. Considering the delete operation can be easily achieved without video content referring, we replace the reference caption with negative captions that have partially misaligned semantics with the video content. Such updated quadruples require models to prioritize removing visually irrelevant content from the reference, which makes the delete more challenging.
Fine-grained attribute-related commands. We construct attribute samples in a “degradation” manner, that is firstly detecting attribute words in the reference caption and then removing them to obtain the edited caption . We utilize Spacy Syntactic Dependency333https://spacy.io/ and Semantic Role Labeling (Shi and Lin, 2019) to achieve noun, verb or modifier attributes detection and removal in while maintaining the fluency of to get . After degradation, we can obtain quadruples for the del, -, attr command. We exchange the reference caption and the edited caption to obtain the opposite add command. The position information can be naturally recorded to support position-related commands.
@@ 6.1. Implementation Details
We implement the small-scale OPA model based on Huggingface Transformers library (Wolf et al., 2020). The default setting is initialized by the . We get video frames using fps=1. For the VATEX-EDIT dataset in English, we adopt BLIP (Li et al., 2022a) ViT-B/16 to extract frame features. The max frame sequence N is set to 20. For the EMMAD-EDIT dataset in Chinese, we initialize our model with the Chinese BART. We adopt CN-CLIP (Yang et al., 2022) ViT-B-16 to extract video frame-level features. The max frame sequence N is set to 30 and the max decoding length is set to 150. For training, we use AdamW (Loshchilov and Hutter, 2017) with a learning rate of 1e-5 and optimize for 20 epochs with a batch size of 20. During inference, we set beam size as 5. The ImgLLM pipeline utilizes the identical frame number N as the OPA model. This pipeline doesn’t involve any training. We choose one in-context learning sample for every command type integrated into the ChatGPT prompt. For VideoChat model, we set frame number N as 8 to fit its default setting. We conduct further instruct-tuning on the official 7B checkpoints with batch size 64.
@@ 6.3. Further Task Analysis
We conduct further ablation studies on the small-scale OPA model to delve into a detailed analysis of the characteristics of the VCE task.
@@ 6.4. Quantitative Results
| Dataset | Model | Control. | Fluency | Vision Align |
| EMMAD-EDIT | Trans. | 2.94 | 3.02 | 3.27 |
| OPA | 3.98 | 3.85 | 3.76 | |
| GT | 4.67 | 4.37 | 4.22 | |
| VATEX-EDIT | Trans. | 4.18 | 4.23 | 3.76 |
| OPA | 4.36 | 4.43 | 3.93 | |
| GT | 4.48 | 4.34 | 4.41 |

### v3

@@ Abstract.
Automatically narrating videos in natural language complying with user requests, i.e. Controllable Video Captioning task, can help people manage massive videos with desired intentions. However, existing works suffer from two shortcomings: 1) the control signal is single-grained which can not satisfy diverse user intentions; 2) the video description is generated in a single round which can not be further edited to meet dynamic needs. In this paper, we propose a novel Video Caption Editing (VCE) task to automatically revise an existing video description guided by multi-grained user requests. Inspired by human writing-revision habits, we design the user command as a pivotal triplet {operation, position, attribute} to cover diverse user needs from coarse-grained to fine-grained. To facilitate the VCE task, we automatically construct an open-domain benchmark dataset named VATEX-EDIT and manually collect an e-commerce dataset called EMMAD-EDIT. We further propose a specialized small-scale model (i.e., OPA) compared with two generalist Large Multi-modal Models to perform an exhaustive analysis of the novel task. For evaluation, we adopt comprehensive metrics considering caption fluency, command-caption consistency, and video-caption alignment. Experiments reveal the task challenges of fine-grained multi-modal semantics understanding and processing. Our datasets, codes, and evaluation tools are available at https://github.com/yaolinli/VCE.
@@ 2. Related Work
Image Caption Editing. Conventional image captioning (Vinyals et al., 2015; Xu et al., 2015; Liu et al., 2016; Stefanini et al., 2022; Li et al., 2016; Liu et al., 2018; Yao et al., 2023; Yang et al., 2023a; Yao et al., 2022) generates the description for images from scratch which may lead to factual mistakes. Sammani and Elsayed (2019) firstly define the image caption editing task that modifies an existing caption (a.k.a. reference caption) conditioned on the image content to obtain more accurate descriptions. Sammani and Melas-Kyriazi (2020) further propose a novel EditNet framework to achieve interactive and adaptive edits. Yuan et al. (2023) design an adaptive text-denoising network to alleviate the semantic gap between input images and reference sentences. The above works all edit image captions implicitly. Wang et al. (2022) propose the explicit image caption editing task to make the modification process more explainable and efficient. In summary, all these works can only correct the wrong content in the reference captions and ignore specific edit intentions of different users. In this paper, we integrate multi-grained user commands into the video description editing process.
Large Multimodal Models. Recent months have witnessed the tremendous success of Large Language Models (LLMs) (Chowdhery et al., 2022; Ouyang et al., 2022; Zeng et al., 2022; Touvron et al., 2023b, a; OpenAI, 2022, 2023a) towards artificial general intelligence. Further, Large Multimodal Models (LMMs) (Dai et al., 2023; Liu et al., 2023; Bai et al., 2023; Ye et al., 2023a; Li et al., 2023a; Zhang et al., 2023; Lin et al., 2023; Ren et al., 2023) endow LLMs with the visual understanding ability by incorporating vision backbones (Sun et al., 2023; Radford et al., 2021; Li et al., 2022b). Existing LMMs primarily bifurcate into two categories: Image Large Language Models (ImgLLMs) (Dai et al., 2023; Liu et al., 2023; Bai et al., 2023; Ye et al., 2023a; Cha et al., 2023) and Video Large Language Models (VidLLMs) (Li et al., 2023a; Zhang et al., 2023; Lin et al., 2023; Luo et al., 2023; Ren et al., 2023). The standard architecture of an LMM comprises a vision backbone for encoding images or videos, a projector (Li et al., 2023b; Cha et al., 2023; Yao et al., 2024) to translate visual embeddings into textual semantic space, and an LLM to process all multimodal contexts. As the VCE task involves capturing video semantics, understanding textual user controls, and enabling text editing, it can serve as a new touchstone for LMMs.
@@ 3.2. User Edit Command
It is not trivial to define flexible edit commands in the VCE task to meet various realistic user needs. We observe that natural language and writing-revision traces are two natural interactive modes. The former can be received from keyboards or speech converters, while the latter conveniently expresses user intentions with the prevalence of tablets and wireless stylus pens. A command representation compatible with the above two signals is important and meaningful. In this paper, we propose a novel command representation in a triplet format {operation, position, attribute}, where operations control the overall description editing, positions specify the editing locations, which could affect the syntax of sentences, and attributes guide the editing operation to control the semantic contents of descriptions.
We define the atomic edit operations as add and delete, considering that the replace editing can be decomposed into the two atomic operations (i.e. first delete then add). Meanwhile, position and attribute in the triplet are optional, therefore, as shown in Table 1, seven specific commands111Note that we omit the command “del, pos, attr, delete attributes at specified positions”, as it can be covered by “del, pos, -, delete description at specified positions”. via different combinations of operation, position, and attribute elements in the triplet can cover multi-grained realistic demands from coarse-grained (global) controls to fine-grained (local) controls. The designed triplet command can be flexibly obtained by processing the inputs from front-end interfaces including natural language and writing-revision traces (details in Appendix E). In the following method and experiments sections, we perform video description editing directly based on the triplet command.
@@ 4.1. VATEX-EDIT Construction
We sample an annotated caption of a video as the reference caption, and the next goal is to construct the command and the related edited caption. In general, we aim to construct related (command, edited caption) samples including: 1) coarse-grained length-control commands referring to the global add or delete edits that result in length changes, and 2) finer-grained attribute-related commands to achieve add or delete attributes.
Coarse-grained length-control commands. For add operation, we directly select a longer caption with remarkable length differences as the edited caption. The delete operation does the exact opposite. Considering the delete operation can be easily achieved without video content referring, we replace the reference caption with negative captions that have partially misaligned semantics with the video content. Such updated quadruples require models to prioritize removing visually irrelevant content from the reference, which makes the delete more challenging.
Fine-grained attribute-related commands. We construct attribute samples in a “degradation” manner, that is firstly detecting attribute words in the reference caption and then removing them to obtain the edited caption . We utilize Spacy Syntactic Dependency222https://spacy.io/ and Semantic Role Labeling (Shi and Lin, 2019) to achieve noun, verb or modifier attributes detection and removal in while maintaining the fluency of to get . After degradation, we can obtain quadruples for the del, -, attr command. We exchange the reference caption and the edited caption to obtain the opposite add command. The position information can be naturally recorded to support position-related commands.
@@ 6.1. Implementation Details
We implement the small-scale OPA model based on Huggingface Transformers library (Wolf et al., 2020). The default setting is initialized by the . We get video frames using fps=1. For the VATEX-EDIT dataset in English, we adopt BLIP (Li et al., 2022a) ViT-B/16 to extract frame features. The max frame sequence N is set to 20. For the EMMAD-EDIT dataset in Chinese, we initialize our model with the Chinese BART. We adopt CN-CLIP (Yang et al., 2022) ViT-B-16 to extract video frame-level features. The max frame sequence N is set to 30 and the max decoding length is set to 150. For training, we use AdamW (Loshchilov and Hutter, 2017) with a learning rate of 1e-5 and optimize for 20 epochs with a batch size of 20. During inference, we set the beam size of generation as 5. The ImgLLM pipeline utilizes the identical frame number N as the OPA model. This pipeline doesn’t involve any training. We choose one in-context learning sample for every command type integrated into the ChatGPT prompt. For VideoChat model, we set frame number N as 8 to fit its default setting. We conduct further instruct-tuning on the official 7B checkpoints with batch size 64.
@@ 6.3. Further Task Analysis
We conduct further ablation studies on the small-scale OPA model to delve into a detailed analysis of the VCE task.
@@
Dataset Model Control. Fluency Vision Align EMMAD-EDIT Trans. 2.94 3.02 3.27 OPA 3.98 3.85 3.76 GT 4.67 4.37 4.22 VATEX-EDIT Trans. 4.18 4.23 3.76 OPA 4.36 4.43 3.93 GT 4.48 4.34 4.41